OpenStudy (anonymous):

What is 15 degrees Celsius (°C) in degrees Fahrenheit (°F)? Use the formula F = 1.8C + 32. A. 20.5 °F B. 37 °F C. 59 °F D. 84.6 °F

OpenStudy (anonymous):

So if it's 15'Celsius, from the formula you know that C=15 So F = (1.8 * C) + 32 F = (1.8 * 15) + 32 we can also write 1.8 as a fraction which is 9/5 so 15 *(9/5) = 27 F = 27 + 32 = ???

OpenStudy (anonymous):

59

OpenStudy (anonymous):

Correct! Good work!